OpenStudy (anonymous):

Evaluate the expression: v ⋅ w Given the vectors: r = <5, -5, -2>; v = <2, -8, -8>; w = <-2, 6, -5>

OpenStudy (zehanz):

This is the dot product of the vectors v and w. The dot product is defined as: if v = (v1,v2,v3) and w=(w1,w2,w3) then v ⋅ w = v1w1+v2w2+v3w3. Just substitute the numbers to see what will come out...

OpenStudy (anonymous):

So the answer would be -12?

OpenStudy (zehanz):

2*-2+-8*6+-8*-5=-4-48+40=-12, yes you are right!

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (zehanz):

yw!